# Self-supervised Learning for Geospatial AI: A Survey

### v1

@@ II-C1 Graph Neural Networks
$h_{v}^{(l)}=\mathcal{F}^{(l)}\left(h_{v}^{(l-1)},\operatorname{AGG}^{(l)}\left(\left\{h_{u}^{(l-1)}\right\}_{u\in\mathcal{N}(v)}\right)\right)$ (3)
where $h_{v}^{(l)}$ indicates the representation of $v$ at layer $l$ and $N(v)$ denotes the neighbors of $v$. $\operatorname{AGG}^{(l)}$ is the message aggregation function at layer $l$, which collects and combines node features, and potentially edge features, from the neighbors, and $\mathcal{F}^{(l)}$ is the function that updates the representation of $v$ based on the aggregated information. For geospatial objects, GNN are frequently utilized to model discrete objects, enabling the capture of complex relationships among them.
II-C2 Sequential Models
@@ II-C3 Pre-trained Models
The evolution of advanced sequential models, especially those based on the Transformer architecture, have marked the milestone in the development of pre-trained language models. One popular paradigm is to adhere to the principles set by BERT [38]. These models [39, 40, 41, 42] leverage large-scale datasets to employ the training objective of masked language modeling (MLM). This process enables the acquisition of rich and transferable representations, which can be fine-tuned for specific tasks with less labeled data. Another paradigm is to employ training with autoregressive language modeling (i.e., next token prediction), exemplified by ChatGPT and its counterparts [43, 44, 45, 46, 47]. These large language models (LLMs) demonstrate strong capabilities in language understanding and reasoning, transforming various tasks into the process of autoregressive language generation. The robust performance of these two paradigms has led to the widespread use of pre-trained language models to encode textual information associated with geospatial data or to adapt their training objectives to develop specialized representations. On the other hand, in scenarios where vision information, such as street view images, is associated with geospatial elements like polylines or polygons, pre-trained visual models are also utilized. These include CNN-based models [48, 49] and recent Transformer-based models [50, 51] trained on large-scale image datasets. When both textual and visual data sources are available, visual-language pre-trained models such as CLIP and its variants [52, 53] are employed to synergize the semantics of the two data modalities, enhancing the interpretability and utility of combined data sources in geospatial applications.
Figure 2: Overview of SSL framework for geospatial data.
@@
In the following sections, for each data type under consideration, our analysis adopts a structured way. We first present the encoding methods for intrinsic attributes that reflect the fundamental characteristics of the data instance. Then we discuss the context information leveraged to enhance the extraction of meaningful insights and knowledge. Finally, we introduce the downstream applications that benefit from the derived geospatial representations. The overview of applying SSL on geospatial data is illustrated in Fig. 2.
III Points
Points are the most fundamental geospatial objects, forming the basic component of polylines and polygons. Each point is associated with a geo-location and associated features, such as text descriptions and categories. Recent studies have focused on learning effective representations for points with SSL techniques, enabling better accuracy in downstream GeoAI applications. As the most representative type of geospatial point data, points-of-interests (POIs) are significant in understanding user mobility and urban functionality. This section will primarily discuss POIs and the corresponding SSL techniques designed for them. The surveyed studies for road networks are listed in Table I
III-A Points-of-Interest
III-A1 Intrinsic Attributes
@@
POIs refer to the semantic locations in location-based services that users might be interested in visiting. Examples of POIs include restaurants, stores, schools, etc. Each POI is associated with geographical coordinates $l$ (i.e., geo-location) and usually a number of features $x=\{name,categories,reviews\}$, including the POI’s name, one or multiple categories, and possibly a set of user reviews.
@@
III-A2 Context Information
Spatial neighborhood. To effectively capture the spatial context of a POI, Skip-gram [57], GNNs [58, 59, 60] and masked language modeling (MLM) [61, 54, 55, 62] have been explored. The skip-gram methods consider the spatial neighbors as the context of a POI, analogous to the surrounding words in Word2vec [63]. They minimize the errors in predicting the spatial context using the target POI embeddings. For instance, Place2vec [57] partitions the spatial neighborhood of a target POI into equal-distance bins and calculates a bin-wise boosting factor to increase the occurrence of popular POI categories in each bin. In this way, POI categories with a higher boosting factor will contribute more to the skip-gram learning objective. In this vein, Huang et al. [64] propose to preserve the hierarchical structure of POI categories in learning POI category embeddings. For example, the categories Japanese Restaurants and Chinese Restaurants should be similar in the embedding space due to their resemblance of functions and the fact that they often share the same generic category, e.g., Food. In this regard, they employ Laplacian eigenmaps as regularization terms to pull together the POI categories that share the same ancestor categories. This method is later extended in [65] and [66] to consider categories with close semantics in different years and the check-in sequences, respectively.
GNN-based methods construct a graph where two POIs are connected by an edge if they are close in spatial. Leveraging the ability of GNNs to acquire the local structure of a graph, these methods encode the spatial neighborhood of a POI into its latent representation. To learn the POI representation in a self-supervised manner, predictive SSL methods have been exploited to predict either the intrinsic attributes of the POI or the relation between POIs. STPA [58] constructs a Delaunay triangulation graph based on the distance between POIs. Each node in the graph is a POI, represented by a one-hot encoding of its category. GNN is applied to acquire a target POI’s representation by aggregating the category information from its neighbors in the Delaunay triangulation graph. Subsequently, a predictive objective is employed to predict the category of the target POI, given its latent representation. For objectives that predict the relations (competitive or complementary) between POIs, the relations between POIs are constructed based on certain heuristic rules. DeepR [59] builds the competitive relations between two POIs if they are close in spatial and frequently co-occur in the same query. DeepR employs a Heterogeneous POI Information Network (HPIN) to represent POI, brands, aspects, and their relations, and a spatial adaptive GNN to acquire the POI representations from neighbors. Given the representations of two POIs, it predicts whether they are in a competitive relation. PRIM [60] constructs competitive relations between different categories of POIs if they frequently co-occur in the same query. Likewise, it constructs complementary relations between POIs in the same category that frequently co-occur in the same query. To predict the relations, PRIM employs the POI representations obtained through a GNN that gives importance to the neighbors based on spatial distance and feature similarity.
Besides, numerous studies leverage the MLM idea applied in pre-trained language models to learn POI representations [61, 55, 54]. The common idea behind MLM-based methods is to construct pseudo sentences from the spatial neighborhoods and then apply MLMs on the pseudo sentences. GeoBERT [61] divides the digital map into grids and proposes two methods to create a pseudo sentence from POIs residing in each grid cell. The first method creates the pseudo sentence by finding the shortest path between the two farthest POIs, passing through all POIs in the same grid cell. The second method returns the ordered sequence of POIs by their distance to the grid center. With the obtained pseudo sentences, GeoBERT treats each POI as a token and adopts the same training mechanism as existing MLMs – predicting a masked POI in the pseudo sentence based on others. SpaBERT [55] creates pseudo sentences by concatenating the names of neighboring POIs. For each POI, SpaBERT finds its neighboring POIs using Geohash and sorts them by their distance to the POI. Subsequently, SpaBERT masks some or all words of a certain POI in a pseudo-sentence and predicts the masked words based on the remaining words in the sentence. In this way, the learned POI representation preserves essential information for generating the words that appear in the nearby POIs, thus capturing the underlying correlations between geo-locations and text. To better leverage the spatial context and support query-POI matching, MGeo [54] designs several geospatial encoders to encode the spatial attributes of a POI, including ID, shape, map position, and relative location to neighboring spatial objects extracted from OpenStreetMap [67]. Subsequently, MGeo employs both predictive and contrastive SSL objectives to train these encoders. The predictive SSL objective minimizes the loss of predicting the masked attribute (e.g., shape) using other attributes. The contrastive objective minimizes the difference between the actual spherical distance and the distance computed by the representations of all pairs of POIs in the same batch. After training the encoders, MGeo fuses the text and spatial information through another MLM that predicts the masked words or spatial attributes of a POI.
Existing studies [68, 69, 70, 71, 72] employ skip-gram or contextual bag-of-word (CBOW), which are two typical predicative objectives, to the check-in sequences. Specifically, C-WARP [68] learns POI representations to predict other POIs within the same check-in sequences. As a result, the co-occurrences of POIs in the same check-in sequences are preserved in the POI representations. Building upon this, CAPE [70] learns POI representations that effectively predict both IDs and texts of other POIs in the same sequences. In this way, the text information in the sequential context is encoded in the learned POI representations. Unlike previous methods that predict all POIs visited within a close time period, DeepMove [71] focuses more on the intent of the trip and considers only the origin and destination POIs of a trip. Utilizing an origin POI, DeepMove predicts other origin POIs with the same destination POI. Consequently, POIs with similar travel intents are projected into a close region in the learned representation space. Hier-CEM [72] extends the check-in sequence context by associating each check-in POI with the ancestor categories. POI2Vec [69] introduces a binary region tree to enhance the modeling of check-in sequences with spatial proximity.
Co-query Context. In digital map services, users can submit text queries to search for POIs. The POIs clicked by a user in the same query are often correlated. Motivated by this, several studies have been dedicated to incorporating the co-query context in learning POI representations. ERNIE-GeoL [62] constructs a heterogeneous graph containing query and POI nodes to capture the correlation between POIs in the same query. POIs are connected by a typed edge if they are clicked successively in the same query or co-locate in the same geographical region. A POI node is connected to its historical query nodes. On the heterogeneous graph, ERNIE-GeoL runs a random walk algorithm to create node sequences, on which it applies two predictive SSL objectives to learn the representations of queries and POIs. Specifically, the first SSL objective minimizes the prediction errors for the masked words of POI or query nodes in the node sequence. The second SSL objective minimizes the error of predicting the geocoding in the discrete global grid system [73]. Besides, to enhance the learning of the text features of POIs, ERNIE-GeoL augments the text attribute of POIs by randomly swapping words with misspelled words or random words. POI relationship prediction methods, including DeepR [59] and SEENet [74], use the POI queries to construct competitive and complementary relations between POIs and employ GNNs to obtain POI representations that preserve the co-query information.
III-A3 Applications
The POI representations acquired from the SSL methods can be applied to various downstream applications. Most POI representations can be used to predict specific POI attributes, such as category [58] and geo-location [54]. Besides, POI presentations can be directly used or used with a subsequent sequential model to predict the next POI visited by a user [68, 69, 70]. The POI representations can also be used to answer POI queries by matching POI representations and the query representation [62]. Furthermore, the presentations of multiple POIs can be used for more complex geospatial data mining tasks. For instance, POI relation prediction methods [59, 60, 74] use the representations of two POIs to predict their competitive or complementary relations. Likewise, the POI representations are pivotal in matching POIs from different databases for entity resolution [76]. In addition, increasing studies have utilized POI embeddings in an aggregated manner for the inference of land use [64] or land use change [65] across years. The aforementioned applications demonstrate that the SSL methods can effectively encode POIs and their contexts, thereby helping to improve the accuracy of various GeoAI tasks.
@@ III-A3 Applications
TABLE I: A summary of the surveyed papers on self-supervised learning for points-of-interest
@@
CatEM [66] Predictive None
Spatial neighbors,
Check-in sequence
Skipgram
Category prediction
@@
C-WARP [68] Predictive None Check-in sequence
@@
Word substitution and masking
@@
SEENet [74] Contrastive Grid cell augmentation
Co-query, Time,
Spatial neighbors
JS divergence
Time-aware POI relation
prediction
IV Polylines

### v2

@@ 2.3.1 Graph Neural Networks
$h_{v}^{(l)}=\mathcal{F}^{(l)}\left(h_{v}^{(l-1)},\operatorname{AGG}^{(l)}\left(\left\{h_{u}^{(l-1)}\right\}_{u\in\mathcal{N}_{v}}\right)\right)$ (3)
where $h_{v}^{(l)}$ indicates the representation of $v$ at layer $l$ and $\mathcal{N}_{v}$ denotes the neighbors of $v$. $\operatorname{AGG}^{(l)}$ is the message aggregation function at layer $l$, which collects and combines node features, and potentially edge features, from the neighbors, and $\mathcal{F}^{(l)}$ is the function that updates the representation of $v$ based on the aggregated information. For geospatial objects, GNNs are frequently utilized to model discrete objects, enabling the capture of complex relationships among them.
2.3.2 Sequential Models
@@ 2.3.3 Pre-trained Models
The evolution of advanced sequential models, especially those based on the Transformer architecture, have marked the milestone in the development of pre-trained language models. One popular paradigm is to adhere to the principles set by BERT [37]. These models [38, 39, 40, 41] leverage large-scale datasets to employ the training objective of masked language modeling (MLM). This process enables the acquisition of rich and transferable representations, which can be fine-tuned for specific tasks with less labeled data. Another paradigm is to employ training with autoregressive language modeling, exemplified by ChatGPT and its counterparts [42, 43, 44, 45, 46]. These large language models (LLMs) demonstrate strong capabilities in language understanding and reasoning, transforming various tasks into the process of autoregressive language generation. The robust performance of these two paradigms has led to the widespread use of pre-trained language models to encode textual information associated with geospatial data or to adapt their training objectives to develop specialized representations. On the other hand, in scenarios where vision information, such as street view images, is associated with geospatial elements like polylines or polygons, pre-trained visual models are also utilized. These include CNN-based models [47, 48] and recent Transformer-based models [49, 50] trained on large-scale image datasets. When both textual and visual data sources are available, visual-language pre-trained models such as CLIP and its variants [51, 52] are employed to synergize the semantics of the two data modalities, enhancing the interpretability and utility of combined data sources in geospatial applications.
Figure 2: Overview of SSL framework for geospatial objects.
(a) Predictive SSL methods
(b) Contrastive SSL methods
Figure 3: Illustration of two types of SSL methods for geospatial objects.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=1em, edge+=darkgray, line width=1pt, s sep=10pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ SSL for Different Geospatial Data Instances, leaf-head, ver [ POIs, middle,text width=8em [ Spatial Neighborhood, middle, text width=10.5em, [Predictive: Place2vec [53], Semantic [54], MT-POI [55], STPA [56],
DeepR [57], PRIM [58], GeoBERT[59], SpaBERT [60]
Predictive&Contrastive: MGeo [61], leaf, text width=38em ] ] [ Check-in Sequence, middle, text width=10.5em [Predictive: SG-CWARP [62], CAPE [63], DeepMove [64], Hier-CEM [65],
POI2Vec [66], TALE [67], LSPSL [68] , leaf, text width=38em ] ] [ Co-query Context, middle, text width=10.5em [Contrastive: MoCo-GA [69] , leaf, text width=38em ] ] [ Multiple Sources, middle, text width=10.5em [Predictive: CatEM [70], ERNIE-GeoL [71]
Contrastive: SEENet [72], DCHL [73], POI-Enhancer [74]
Predictive&Contrastive: STKG-PLM [75], AGCL [76] , leaf, text width=38em] ] ] [ Road Networks, middle,text width=8em [ Intrinsic Attributes, middle, text width=10.5em, [ Predictive: R2Vec [77], IRN2Vec [78], RFN [79], HyperRoad [80], leaf, text width=38em ] ] [ Spatial Attributes, middle, text width=10.5em, [Predictive: RN2Vec [81]   Contrastive: SARN [82], leaf, text width=38em ] ] [ Trajectories, middle, text width=10.5em [Predictive: HNRN [83], Toast [84], DyToast [85], TrajRNE [86]
Contrastive: JCLRNT [87]   Predictive&Contrastive: TRACK [88], leaf, text width=38em ] ] [ Street View Images, middle, text width=10.5em [Contrastive: Garner [89]   Predictive&Contrastive: USPM [90], leaf, text width=38em ] ] ] [ Trajectory, middle,text width=8em [ Intrinsic Attributes, middle, text width=10.5em, [Predictive: TCDRL [91], t2vec [92], GM-VSAE [93], Geo-Tokenizer [94],
AdvTraj2vec [95], E2DTC [96], STPT [97], ST-t2vec [98], DeepTEA [99], UniTraj [100]
Contrastive: CL-Tsim [101], TrajCL [102], KGTS [103]
Predictive&Contrastive: RSTS [104], leaf, text width=38em ] ] [ Road Networks, middle, text width=10.5em, [Predictive: CSSRNN [105], Traj2Vec [106], PT2Vec [107], JGRM [108]
Contrastive: ST2Vec [109], GRLSTM [110], PIM [111], MMTEC [112]
Predictive&Contrastive: HMTRL [113], START [114], LightPath [115],
GREEN [116], leaf, text width=38em ] ] [ Semantic Information, middle, text width=10.5em [Predictive: CTLE [117], At2vec [118]
Contrastive: SML-LP [119], CLMTR [120], CACSR [121] , leaf, text width=38em ] ] [ Multiple Sources, middle, text width=10.5em [Predictive: RED [122]   Contrastive: PTrajM [123]
Predictive&Contrastive: MVTraj [124], MM-Path [125], leaf, text width=38em ] ] ] [ Urban Regions, middle,text width=8em [ Intrinsic Attributes, middle, text width=10.5em, [Contrastive: Triplet [126], UrbanVLP [127], MMGR [128]
Predictive&Contrastive: UrbanCLIP [129], leaf, text width=38em ] ] [ POIs/ Knowledge
    Graphs, middle, text width=10.5em, [Contrastive: HGI [130]   Predictive&Contrastive: SKRL4RS [131], leaf, text width=38em ] ] [ Human Trajectories, middle, text width=10.5em [Predictive: HDGE [132], ZE-Mob [133], GMEL [134], MGFN [135], leaf, text width=38em ] ] [ POIs + Human
Trajectories (+ others), middle, text width=10.5em [Predictive: MP-VN [136], DLCL [137], CGAL [138], HAFusion [139]
MVURE [140], Region2Vec[141], HUGAT[142]
Contrastive: ReMVC [143]
Predictive&Contrastive: HREP [144], ReCP [145], leaf, text width=38em ] ] [ Imagery Data, middle, text width=10.5em [Predictive: RegionEncoder [146]
Contrastive: Tile2Vec [147], Urban2Vec [148], M3G [149], Xi et al [150],
GeoHG [151], MuseCL [152]
Predictive&Contrastive: ReFound [153], leaf, text width=38em ] ] [ OpenStreetMap, middle, text width=10.5em [ Contrastive: RegionDCL [154], leaf, text width=38em ] ] ] ]
Figure 4: Taxonomy of SSL methods for geospatial data instances discussed in the paper. It is categorized by the type of context information leveraged to for encoding each data instance.
@@
It is important to note that these representative models are not mutually exclusive, and they can be employed concurrently to encode geospatial objects through model combination or to incorporate diverse contextual information. In the following sections, we adopt a structured analytical approach for each geospatial data type under consideration. We first present the encoding methods for intrinsic attributes that reflect the fundamental characteristics of the data instance. Then we discuss methods to integrate and fuse heterogeneous context information to enhance the extraction of meaningful insights and knowledge for the studied geospatial object. Finally, we introduce the downstream applications that benefit from the derived geospatial representations.
2.4 Overview and Taxonomy
Figure 2 provides an overview of the Self-Supervised Learning (SSL) framework for geospatial objects, illustrating the pipeline from raw data to downstream applications. At the foundation, geospatial objects are categorized into three primary geometric data types–points, polylines, and polygons–each corresponding to different data instances such as POIs, road segments, trajectories, and urban regions. These objects and their context information and augmented variations serve as input to the SSL framework. Within this framework, predictive methods focus on recovering missing or masked attributes, while contrastive methods aim to learn representations by distinguishing between similar and dissimilar instances. These techniques are guided by specific pretext tasks and objective functions, such as the cross-entropy for attribute classification, or the InfoNCE loss for contrastive learning. Lightweight task-specific models then utilize the learned representations to support various geospatial applications, such as POI classification, traffic prediction, anomaly detection, socio-economic inference, etc.
The taxonomy of SSL methods discussed in the paper is presented in Figure 4. It is organized according to data instance and the contextual information exploited for representation learning. The taxonomy comprises four types of geospatial data instances, each grouped according to the specific context information they utilize. We classify SSL methods within each branch into predictive, contrastive, or hybrid (predictive & contrastive) methods. To enhance conceptual clarity, Figure 3 provides a schematic illustration for predictive and contrastive SSL methods, as introduced in Section 2.2, applied across different geospatial objects. This illustration outlines the core modeling workflows and highlights the general design principles underlying each SSL paradigm. Building on these paradigms, each geospatial data type incorporates tailored designs for the constituent modules and data sources within its respective learning pipeline. In the following sections, we present a detailed discussion of the methods under each branch of the taxonomy.
3 Points
Points are the most fundamental geospatial objects, forming the basic component of polylines and polygons. Each point is associated with a geo-location and associated features, such as text descriptions and categories. Examples of point data include location-based point sensors, road network intersections, and points-of-interest (POIs). While extensive research has been conducted on modeling point data for forecasting tasks related to traffic, weather, and environment [155, 156], most studies adopt an end-to-end training method to improve forecasting accuracy, with no emphasis on representation learning. Recent research on self-supervised representation learning techniques for point data focuses primarily on POI data due to their rich semantic information. Consequently, this survey focuses on POI data, which plays a crucial role in understanding user mobility and urban functionality. The surveyed studies for POI data are listed in Table 1.
3.1 Points of Interest
3.1.1 Intrinsic Attributes
@@
POIs refer to the semantic locations in location-based services that users might want to visit. Examples of POIs include restaurants, stores, schools, etc. Each POI is associated with geographical coordinates $l$ (i.e., geo-location) and usually a number of features $x=\{name,categories,reviews\}$, including the POI’s name, one or multiple categories, and possibly a set of user reviews.
@@
Table 1: A summary of the surveyed papers on self-supervised learning for points-of-interest.
@@
SG-CWARP [62] Predictive None Check-in sequence
@@
TALE [67] Predictive Binary temporal tree construction Check-in sequence CBOW POI prediction
LSPSL [68] Predictive POI attribute masking Check-in sequence Cross-Entropy Masked POI attribute prediction
MoCo-GA [69] Contrastive Samples from previous training steps Co-query InfoNCE
Location-location,
Text-location,
Text-text contrast
SEENet [72] Contrastive Grid cell augmentation
Temporal context,
Spatial neighbors,
Co-query,
check-in sequence
JS divergence
Time-aware POI relation
prediction
CatEM [70] Predictive None
Spatial neighbors,
Check-in sequence
Skipgram
Category prediction
DCHL [73] Contrastive Multi-view hypergraph construction
Spatial neighbors
Check-in sequence
InfoNCE POI-POI contrast
STKG-PLM [75]
Predictive,
Contrastive
Spatio-temporal knowledge
graph construction
Spatial neighbors
Check-in sequence
InfoNCE
Cross-Entropy
POI-POI contrast
POI prediction
AGCL [76]
Predictive,
Contrastive
Spatio-temporal knowledge
graph, co-occurrence frequency
graph construction
Spatial neighbors
Check-in sequence
InfoNCE
Cross-Entropy
POI-POI constrast
POI prediction
POI-Enhancer [74] Contrastive
Geography view,
Sequence-time view,
Functional view
construction
Spatial neighbors
Check-in sequence
InfoNCE POI-POI contrast
@@
Word substitution
and masking
@@
3.1.2 Context Information
Spatial Neighborhood. To effectively capture the spatial context of a POI, Skip-gram [53], GNNs [56, 57, 58] and masked language modeling (MLM) [59, 61, 60, 71] have been explored. The skip-gram methods consider the spatial neighbors as the context of a POI, analogous to the surrounding words in Word2vec  [158]. They minimize the errors in predicting the spatial context using the target POI embeddings. For instance, Place2vec [53] partitions the spatial neighborhood of a target POI into equal-distance bins and calculates a bin-wise boosting factor to increase the occurrence of popular POI categories in each bin. In this way, POI categories with a higher boosting factor will contribute more to the skip-gram learning objective. In this vein, Huang et al. [54] propose to preserve the hierarchical structure of POI categories in learning POI category embeddings. For example, the categories Japanese Restaurants and Chinese Restaurants should be similar in the embedding space due to their resemblance of functions and the fact that they often share the same generic category, e.g., Food. They employ Laplacian eigenmaps as regularization terms to pull together the POI categories that share the same ancestor categories. This method is later extended in [55] and [70] to consider categories with close semantics in different years and the check-in sequences, respectively.
GNN-based methods construct a graph where two POIs are connected by an edge if they are close in spatial. Leveraging the ability of GNNs to acquire the local structure of a graph, these methods encode the spatial neighborhood of a POI into its latent representation. Predictive SSL methods have been proposed to predict either the intrinsic attributes of the POI or the relation between POIs. STPA [56] constructs a Delaunay triangulation graph based on the distance between POIs. Each node in the graph is a POI, represented by a one-hot encoding of its category. GNN is applied to acquire a target POI’s representation by aggregating the category information from its neighbors in the Delaunay triangulation graph. Subsequently, a predictive objective is employed to predict the category of the target POI, given its latent representation. For objectives that predict the relations (competitive or complementary) between POIs, the relations between POIs are constructed based on certain heuristic rules. DeepR [57] builds the competitive relations between two POIs if they are spatially close and frequently co-occur in the same query. DeepR employs a Heterogeneous POI Information Network (HPIN) to represent POI, brands, aspects, and their relations, and a spatial adaptive GNN to acquire the POI representations from neighbors. Given the representations of two POIs, it predicts whether they are in a competitive relation. PRIM [58] constructs competitive relations between different categories of POIs if they frequently co-occur in the same query. Likewise, it constructs complementary relations between POIs in the same category that frequently co-occur in the same query. To predict the relations, PRIM employs the POI representations obtained through a GNN that gives importance to the neighbors based on spatial distance and feature similarity.
Besides, numerous studies leverage the MLM idea applied in pre-trained language models to learn POI representations [59, 60, 61]. The common idea behind MLM-based methods is to construct pseudo sentences from the spatial neighborhoods and then apply MLMs to the pseudo sentences. GeoBERT [59] divides the digital map into grids and proposes two methods to create a pseudo sentence from POIs residing in each grid cell. The first method creates the pseudo sentence by finding the shortest path between the two farthest POIs, passing through all POIs in the same grid cell. The second method returns the ordered sequence of POIs by their distance to the grid center. With the obtained pseudo sentences, GeoBERT treats each POI as a token and adopts the same training mechanism as existing MLMs – predicting a masked POI in the pseudo sentence based on others. SpaBERT [60] creates pseudo sentences by concatenating the names of neighboring POIs. For each POI, SpaBERT finds its neighboring POIs using Geohash and sorts them by their distance to the POI. Subsequently, SpaBERT masks some or all words of a certain POI in a pseudo-sentence and predicts the masked words based on the remaining words in the sentence. In this way, the learned POI representation preserves essential information for generating the words that appear in the nearby POIs, thus capturing the underlying correlations between geo-locations and text. To better leverage the spatial context and support query-POI matching, MGeo [61] designs several geospatial encoders to encode the spatial attributes of a POI, including ID, shape, map position, and relative location to neighboring spatial objects extracted from OpenStreetMap [159]. Subsequently, MGeo employs both predictive and contrastive SSL objectives to train these encoders. The predictive SSL objective minimizes the loss of predicting the masked attribute (e.g., shape) using other attributes. The contrastive objective minimizes the difference between the actual spherical distance and the distance computed by the representations of all pairs of POIs in the same batch. After training the encoders, MGeo fuses the text and spatial information through another MLM that predicts the masked words or spatial attributes of a POI.
Most existing studies [62, 66, 63, 64, 65] employ skip-gram, contextual bag-of-word (CBOW), which are three typical predicative objectives, to the check-in sequences. Specifically, SG-CWARP [62] learns POI representations to predict other POIs within the same check-in sequences. As a result, the co-occurren ces of POIs in the same check-in sequences are preserved in the POI representations. Building upon this, CAPE [63] learns POI representations that effectively predict IDs and texts of other POIs in the same sequences. In this way, the text information in the sequential context is encoded in the learned POI representations. Unlike previous methods that predict all POIs visited within a close time period, DeepMove [64] focuses more on the intent of the trip and considers only the origin and destination POIs of a trip. Utilizing an origin POI, DeepMove predicts other origin POIs with the same destination POI. Consequently, POIs with similar travel intents are projected into a close region in the learned representation space. Apart from predictive methods that employ the skip-gram objective, SSL methods with the CBOW objective have been proposed to predict a check-in POI based on its context. These methods often focus on the modeling of spatio-temporal context in the check-in sequences. Hier-CEM [65] extends the check-in sequence context by associating each check-in POI with the ancestor categories. POI2Vec [66] introduces a binary region tree to enhance the modeling of check-in sequence context with spatial proximity. TALE [67] extends the binary region tree in POI2Vec with time information, producing a temporal tree that splits the check-ins into different time slots. In this way, TALE captures both spatial and temporal proximity in the context of a check-in.
Motivated by Masked Language Models (MLM), some recent studies propose to learn POI representation by predicting the masked check-in in a sequence. LSPSL [68] obtains a check-in’s representation by a self-attention network, aggregating information from its attributes (e.g., location). The attribute embeddings are then pre-trained by employing each check-in’s representation to predict the attributes of other check-in records in the same trajectory.
Co-query Context. In digital map services, users can submit text queries to search for POIs. The POIs clicked by a user in the same query are often correlated. Motivated by this, several studies have been dedicated to incorporating the co-query context in learning POI representations. ERNIE-GeoL [71] constructs a heterogeneous graph containing query and POI nodes to capture the correlation between POIs in the same query. POIs are connected by a typed edge if they are clicked successively in the same query or co-locate in the same geographical region. A POI node is connected to its historical query nodes. On the heterogeneous graph, ERNIE-GeoL runs a random walk algorithm to create node sequences, on which it applies two predictive SSL objectives to learn the representations of queries and POIs. Specifically, the first SSL objective minimizes the prediction errors for the masked words of POI or query nodes in the node sequence. The second SSL objective minimizes the error of predicting the geocoding in the discrete global grid system [160]. Besides, to enhance the learning of the text features of POIs, ERNIE-GeoL augments the text attribute of POIs by randomly swapping words with misspelled words or random words. POI relationship prediction methods, such as SEENet [72], use the POI queries to construct competitive and complementary relations between POIs and employ GNNs to obtain POI representations that preserve the co-query information. Unlike the previous predictive SSL methods, MoCo-GA [69] employs momentum contrastive learning to ensure the current POI embeddings are closer to the query embeddings than those obtained in the previous training steps.
Multiple Context Sources. Since POIs are often located within complex spatiotemporal contexts, extensive research highlights the value of incorporating multiple contextual sources when learning POI representations. CatEM [70] represents the category co-occurrence in check-in sequences by a Point-wise Mutual Information (PMI) matrix and captures the spatial neighborhood by a category proximity matrix. CatEM learns POI representations by simultaneously optimizing the matrix factorization loss on the PMI matrix and minimizing the distance between embeddings of nearby POIs. In this way, CatEM incorporates the spatial neighborhood context and the check-in sequence context. DCHL [73] constructs hypergraphs from three different views, including geographical view, collaborative view, and transitional view, to incorporate the spatial neighborhood and sequential information of check-in sequences at the same time. To learn POI representations, DCHL employs a cross-view contrastive learning technique that takes the same POI from the different views as positive pairs while different POIs as negative pairs. STKG-PLM [75] models the POI data with a spatio-temporal knowledge graph, defining spatial neighborhoods and check-in context by pre-defined relations. It employs a knowledge graph encoder to obtain POI representations and the InfoNCE [12] objective for minimizing the gap between POI representations checked in by the same user while maximizing those checked in by different users. Similar to STKG-PLM, AGCL [76] constructs a spatio-temporal knowledge graph to capture the spatial neighborhood and transitional relations between POIs. Different from STKG-PLM, it constructs a co-occurrence frequency graph to incorporate the frequency of all users who visit one POI after another. To learn POI representations, AGCL employs both intra-graph and inter-graph contrastive learning objectives. The intra-graph contrastive objective ensures that POIs with high correlations within and across graphs are positioned closely in the embedding space. POI-Enhancer [74] proposes a POI embedding model to extract semantic information of POIs from large language models and employs a multi-view contrastive learning method to learn POI representations. It creates a geography view, a sequence-time view, and a functional view of POIs to incorporate the spatial, sequential, and semantic context of POIs. POIs that are located in a close region are considered positive samples in the geography view, and those that share the same category and visit patterns are considered positive samples in the functional view. In the sequence-time view, POIs visited in a close time period are regarded as positive samples. With these three views, POI-Enhancer learns a comprehensive POI representation by optimizing the InfoNCE loss. While most studies incorporate the spatial context and the sequential context of POIs, some research has been dedicated to exploring other combinations of context sources. For instance, ERNIE-GeoL [71] incorporates multiple context sources by defining a heterogeneous graph with three types of edges, i.e., Origin-to-Destination, POI co-location, and POI co-query. SEENet [72] constructs a dynamic graph based on the spatial proximity and the temporal relation between POIs to incorporate spatiotemporal, sequential, and co-query contexts.
3.1.3 Applications
The POI representations acquired from the SSL methods can be applied to various downstream applications. Most POI representations can be used to predict specific POI attributes, such as category [56] and geo-location [61]. Besides, POI presentations can be directly used or used with a subsequent sequential model to predict the next POI visited by a user [62, 66, 63]. The POI representations can also be used to answer POI queries by matching POI representations and the query representation [71]. Furthermore, the representations of multiple POIs can be used for more complex geospatial data mining tasks. For instance, POI relation prediction methods [57, 58, 72] use the representations of two POIs to predict their competitive or complementary relations. Likewise, the POI representations are pivotal in matching POIs from different databases for entity resolution [162]. In addition, increasing studies have utilized POI embeddings in an aggregated manner for the inference of land use [54] or land use change [55] across years. The aforementioned applications demonstrate that the SSL methods can effectively encode POIs and their contexts, thereby helping to improve the accuracy of various GeoAI tasks.
4 Polylines